[figure: 9.7 cm]

Kondo-Lattice Model: Application to Thin Ferromagnetic Local-Moment Films

W. Nolting

Lehrstuhl Festkörpertheorie, Institut für Physik, Humboldt-Universität zu Berlin, Invalidenstr. 110, D-10115 Berlin

We study correlation effects and temperature dependencies in the electronic structure of thin ferromagnetic local-moment films. In a first step the Kondo-lattice model is investigated as a candidate for a proper representation of local-moment ferromagnets. Magnetic and electronic key-quantities as the Curie-temperature and the quasiparticle density of states are derived with previously tested many-body procedures. It is shown that the magnetic properties can be interpreted exclusively in terms of the temperaturedependent electronic quasiparticle structure. An extended RKKY theory leads to effective Heisenberg exchange integrals, which turn out to be functionals of the conduction electron selfenergy, getting therewith a remarkable temperature and band occupation dependence. In a second step the model studies are combined with tight binding-LMTO bandstructure calculations in order to get for real ferromagnetic films quasiparticle densities of states and quasiparticle bandstructures. The proposed method avoids the double-counting of relevant interactions and takes into account the correct symmetry of the atomic orbitals. Special results are given for thin ferromagnetic EuO (100) films. The Curie temperature TC of the EuO film turns out to be strongly thickness-dependent, starting from a very low value (≃ 15 K) for the monolayer and reaching the bulk value at about 30 layers. For a 20-layer film we predict the existence of a surface state, the temperature-behaviour of which can lead to a surface halfmetal-insulator transition.

[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : oral

Corresponding author : W. Nolting

Address for correspondence : Institut für Physik Humboldt-Universität zu Berlin Invalidenstr. 110 D-10115 Berlin

Email address : wolfgang.nolting@physik.hu-berlin.de